# Advancing Compositional Awareness in CLIP with Efficient Fine-Tuning

### v1

@@ 1 Introduction
While SugarCrepe [17] mainly tests whether a model can distinguish between lexically similar but semantically different texts (: “The dog is chasing a cat” vs. : “The cat is chasing a dog”), it has been pointed out in SugarCrepe++ [10] that this is not sufficient for true understanding. Given an image of a dog chasing a cat, models should attach higher similarity of the image to both and a semantically equivalent but lexically dissimilar text (“The cat is being chased by a dog”) than to the wrong text, . Surprisingly, it turns out that improving compositionality on SugarCrepe++ is non-trivial. As can be seen in Table 1, previous works perform very well on SugarCrepe (e.g., DAC [8]), but their performance on SugarCrepe++ is even worse than the pre-trained model.
@@
In this paper, we introduce Compositionally-aware Learning in CLIP (CLIC). This technique fine-tunes only the text-encoder in CLIP models by leveraging already available high-quality captioned datasets like PixelProse [44] or common text-image datasets like Laion [41], which we recaption using CogVLM [47]. Using a novel technique of combining images and captions allows us to generate positives as well as hard-negatives with minimal additional overhead, e.g., we do not require an LLM for generating hard-negatives [8], nor do we need to generate synthetic images [37]. In Figure 1, we show that CLIC consistently improves both on SugarCrepe++ and on image-to-text and text-to-image retrieval benchmarks compared to the pre-trained CLIP-ViT-B/32 model [40], whereas existing methods do not achieve consistent gains. To summarize, our contributions are as follows.
@@ 2 Methods and benchmarks for improving compositionality
Recent studies on VLMs have revealed significant limitations in their compositional abilities [49, 17, 46, 32] via specially designed benchmarks. Throughout these works, several types of image-text models have been shown to lack compositional reasoning, ranging from those using just contrastive loss (CLIP [40], LaCLIP [11], SigLiP [50]) to the ones using different losses or architectures like CoCa [48], FLAVA [42], CLIPA [29], BLIP [15]. In this section, we introduce our notation for standard contrastive learning and provide details on methods to improve compositionality and the associated benchmarks.
@@ 2.2 Compositionality enhancing methods
Dense and Aligned Captions (DAC) [8] uses the same method to create hard negatives, but first generates higher quality captions by passing the image through BLIP-2 [23]. In DAC-LLM, the enhanced captions are used as input to GPT-Neo-2.7B [13] for detailed and dense captions, and in DAC-SAM, segments of the original image, created using SAM [19], are captioned via BLIP-2. The model is then fine-tuned using a combination of (Eq. 1), (Eq. 3) and MIL-loss adapted from [34], which uses multiple negatives for each image.
TripletCLIP [37] uses rewritten captions from LaCLIP [11] and utilizes an LLM to generate hard-negatives by changing concepts such as objects, attributes, and relationships that change the meaning of the original caption while slightly distorting it. Afterward, a synthetic image is generated for the hard-negative captions using a text-to-image diffusion model. Finally, training is done by using a sum of image-text and text-image based from Eq. 2.
@@ 2.3 Compositionality benchmarks
Recently, SugarCrepe++ [10] extended this idea by introducing , an additional caption that is semantically similar to , but lexically different from both and . To answer correctly, the model needs to assign a higher score to and compared to the score assigned to . This makes SugarCrepe++ harder and more meaningful than SugarCrepe, as the model can no longer rely solely on lexical differences but has to instead understand the semantic differences between the captions. Therefore, we predominantly focus on SugarCrepe++ in this work.
@@ 3 Compositionally-aware learning in CLIP
Large-scale web-crawled datasets are known to be noisy, with many captions not reflecting their corresponding images [25, 3, 51]. This can be detrimental to compositional learning, which relies on minor modifications that alter semantic meaning. To improve the compositionality, we opt for captions that better align with the images and describe the images in detail. For this, we use approximately sample subsets taken from different data sources, like, RedCaps and CC12M from PixelProse [44], and our CogVLM [47] recaptioned Laion. Details about the datasets can be found in Appendix B. We train several versions of CLIC using each dataset to demonstrate the generalizability of our approach across data sources. Additionally, since some baselines are trained using MS-COCO, for a fair comparison, we train another version on this dataset, the details of which are in Section 4.
@@ 3.1 Generation of positives and hard-negatives: concatenation of images and captions
To overcome the “bag of words" limitation, we aim for the model to distinguish between captions that are lexically similar but semantically different. In order to achieve this, we construct hard negatives that retain the same words in the caption but no longer correctly describe the image. For efficiency and diversity, we create these examples by concatenating pairs of images and swapping words across their two captions, ensuring that the resulting caption does not represent the concatenated images. To help the model learn that different lexical or syntactic sentences can represent the same image, we create multiple positives. For this, we use dense captions that provide varied descriptions of each image. The concatenation enables the number of combinations to grow quadratically with the dataset size, allowing CLIC to generalize better than other methods, see discussion in Section B.3. We detail our data preparation strategy based on concatenating random image-caption pairs in the following section. The process is illustrated in Figure 2 and described in Algorithm 1 in Appendix C. See Section B.4 for a discussion on a non-random approach.
@@
: We then sample (without repetition) two additional sentences from each caption and concatenate them in a random order , and for random indices . This step reinforces the model’s ability to distinguish incorrect captions from correct ones, even when the incorrect caption presents a general description of the image while the correct caption describes only a specific part. This is different from previous works, such as DAC [8], where multiple positives exist, but each caption faces its own hard-negative. An ablation on the number of positives can be found in Table 13.
: Next, we leverage the spaCy package [16] to generate hard-negatives. We decompose the two first sentences into individual words and identify their respective linguistic categories using spaCy. A table detailing these categories, along with examples, is provided in Table 5 in Appendix B. To construct hard-negatives, we randomly select a linguistic category that is common to both sentences and choose two words from that category – one from each image. If no common category exists, we randomly select a word from each image and swap them. Certain categories, such as punctuation, are excluded from this process (see Appendix B.2 for details). To prevent over-fitting to specific benchmarks, we do not impose constraints on the selection of particular categories or word positions. As a result, the generated sentence no longer accurately represents the concatenated image unless, by chance, the swapped words share the same meaning.
@@ 3.2 Training of CLIC
Hard-negative loss: We compute a hard-negative loss between each of the positive captions and the negative caption of the same image, similar to Eq. 3,
@@
where , , and are hyperparameters controlling the contribution of each loss term. To prevent the model from deviating too much from the pretrained CLIP when training only on concatenated images, we employ the standard clip training every second iteration, using a single image, the first sentence, and the standard clip loss as in Eq. 1.
@@ 3.3 Differences between CLIC and other methods
Other approaches for generating hard text negatives differ from ours in computational cost and methodology. In DAC [8] and SVLC [9], hard-negatives tailored explicitly for specific compositionality tasks (e.g., replacing specific categories like colors, actions, etc.) are either found through look-up tables or created with computationally expensive LLM-based generation. Similarly, TripletCLIP uses an LLM to generate hard-negatives; for details, see Section 2.2. Closest to our work is NegCLIP, which also uses word swapping. However, NegCLIP focuses on specific word classes such as adjectives, adverbs, verbs, and nouns as described in Appendix D. We allow a swap of any two random words as long as they are from the same class, e.g., prepositions, numbers, and pronouns, as seen in Section B.2. Our approach offers two main advantages: cheaper computational cost compared to LLM-based generation and a design that doesn’t target any specific benchmark, reducing the risk of over-fitting on existing evaluation benchmarks. Other methods also add hard-negative images, further increasing overall computational complexity. NegCLIP and CoN-CLIP search for hard images to add to the batch, and TripletCLIP generates hard images using a diffusion model. In addition, none of the methods use a combination of image pairs and corresponding text descriptions as in our method.
@@ 4 Experiments and evaluation
Here, we describe the experimental details and the main results of our work.
@@ 4.1 Experimental setup
Dataset.
For CLIC, we fine-tune models either with CogVLM [47] recaptioned Laion images, or PixelProse [44] recaptioned images from RedCaps and CC12M (see Appendix B for details).
We stress that this approach keeps our training data independent of MS-COCO, which is the basis for benchmarks like SugarCrepe, SugarCrepe++, and retrieval evaluations. By that, we
ensure CLIC always functions as a zero-shot model.
However, to fairly compare with the ViT-B/32 model of NegCLIP, which is trained on MS-COCO, we also fine-tune one model on the MS-COCO [27] dataset.
In our experiments, we denote each model by the respective dataset it was trained on, i.e., CLIC-COCO, CLIC-LAION, CLIC-RedCaps, or CLIC-CC12M.
Models.
To be comparable with previous compositionality enhancing methods, we base the majority of our experiments and the ablations on ViT-B/32 models pre-trained by OpenAI [40].
However, to highlight the adaptability and scalability of our approach, we also fine-tune other commonly used pre-trained architectures, such as ViT-B/16 and ViT-L/14 from [40]. Up to our knowledge, among methods improving compositionality, only CoN-CLIP [43] provides checkpoints for these architectures.
Additionally, we fine-tune more recent ViT-L equivalent VLMs, including CLIPS and CLIPA to show that our method can enhance compositionality even in models that already possess compositional abilities and are pre-trained on varying datasets such as Re-DataComp (CLIPS), WIT (CLIP) and DataComp (CLIPA). More model specific details can be found in Appendix D.
Training details.
In all of our experiments for CLIC, we keep the vision encoder frozen and fine-tune only the text encoder. An ablation showing that this works better than training the entire model can be found
in Table 12 in the appendix.
We find it really intriguing that the main improvement in terms of compositionality can be achieved by just fine-tuning the text encoder.
We defer other training details like LR, loss parameters in Eq. 7, etc, and a discussion on our NegCLIP baselines (denoted everywhere with a †), used whenever there are no available checkpoints to Appendix D.
Evaluation.
We evaluated the models on the compositionality benchmarks SugarCrepe++ [10], WinoGround [46], and SugarCrepe [17]. Additionally, we assess the effects of fine-tuning on downstream tasks such as classification on ImageNet [5] and on ZS-10, which computes the average score across ten standard classification datasets, detailed in Section D.2. In addition, we report image () and text retrieval () Recall@5 scores on MS-COCO (Val2017) [27] and Flickr30k [38]. We show Recall@1 numbers for retrieval and detailed SugarCrepe and SugarCrepe++ numbers for all the models in Appendix E.
@@ 4.2 Comparing to other methods
Given that most publicly available checkpoints of compositionality methods use ViT-B/32 based CLIP, we first focus on this architecture.
From Table 1, we observe that most methods perform worse than the pre-trained CLIP on SugarCrepe++. Notably, only NegCLIP, CoN-CLIP, and the CLIC models improve or are on par with the pre-trained baseline. Among these, CLIC-COCO is better than NegCLIP and CoN-CLIP (all three were trained on MS-COCO), while CLIC-RedCaps has lower TOT with a higher ITT (10.2% gain on pre-trained model across replace and swap). We also evaluate all models on the WinoGround benchmark, where, apart from CLIC, no other method improves on the pre-trained model, highlighting the effectiveness of CLIC. Moreover, CLIC improves in compositionality regardless of the fine-tuning dataset, showing that CLIC works for different sources of captions and images used for fine-tuning.
An interesting finding in Table 1 is that many of the models presented in the table perform well on ITT in SugarCrepe, but struggle for ITT in SugarCrepe++.
We remind the reader that ITT in SugarCrepe++ is an extended version of ITT from SugarCrepe. It additionally
includes lexically different but semantically similar positives, which the model should also favour over the negatives.
This stark difference in performance for methods like DAC (has SugarCrepe replace ITT of 89.4% but SugarCrepe++ ITT of only 53.7%) and SVLC suggests these models might not be truly learning compositional relations. More specifically, this suggests that although the models manage to detect lexical changes from the original text, they struggle when facing a lexically different text with the same meaning. The probable cause of this could be
the over reliance of these methods on using
very specific hard-negatives tuned for
previous benchmarks, see discussion in Section D.4.
Downstream tasks. We evaluate the models on standard retrieval benchmarks (MS-COCO and Flickr30k) and on classification tasks (ImageNet and ZS-10), as detailed in Table 2. By looking at both these tables, one sees that TripletCLIP performs well on SugarCrepe++ and WinoGround, but is found lacking in downstream tasks performance on ImageNet, ZS-10. We see that our CLIC models exhibit lower performance degradation in classification tasks compared to most models. In addition, CLIC-CC12M achieves the highest text retrieval performance among the models not trained on MS-COCO while also improving on the pre-trained model for image retrieval.
@@ 4.3 Ablation Study for CLIC
In Table 10 in Appendix E we ablate the components of CLIC in steps (C1-C5) and present the analogous non-concatenated, single image version of CLIC. The CLIC baseline (C1) is fine-tuning on the concatenated images with caption and in (C2) we add our single hard negative. These two steps are sufficient to improve on SugarCrepe but not on SugarCrepe++. This suggests that the hard-negative distractor helps the model to detect lexical changes, but not to distinguish between semantically similar sentences and ones with different meaning. Adding multiple positives (C3) and the uni-modal loss (C4) improves significantly on SugarCrepe++ while performance on SugarCrepe slightly reduces. Thus adding diverse positives and enforcing invariance of semantically identical captions lets the model identify semantically similar captions while being sensitive to lexically similar but semantically dissimilar captions. In the last step (C5) we add the iterates with single images and standard contrastive loss which further improves compositionality as well as downstream task performance by eliminating the potential bias of the model due to the concatenated images. The non-concatenated, single image variant of CLIC shows similar but smaller improvements in compositionality while being significantly worse in retrieval.

### v2

@@ 1 Introduction
While SugarCrepe [18] mainly tests whether a model can distinguish between lexically similar but semantically different texts (: “The dog is chasing a cat” vs. : “The cat is chasing a dog”), it has been pointed out in SugarCrepe++ [11] that this is not sufficient for true understanding. Given an image of a dog chasing a cat, models should attach higher similarity of the image to both and a semantically equivalent but lexically dissimilar text (“The cat is being chased by a dog”) than to the wrong text, . Surprisingly, it turns out that improving compositionality on SugarCrepe++ is non-trivial. As can be seen in Table˜1, previous works perform very well on SugarCrepe (e.g., DAC [9]), but their performance on SugarCrepe++ is even worse than the pre-trained model.
@@
In this paper, we introduce Compositionally-aware Learning in CLIP (CLIC). This technique fine-tunes CLIP models by leveraging already available high-quality captioned datasets like PixelProse [50] or common text-image datasets like Laion [46], which we recaption using CogVLM [53]. Using a novel technique of combining images and captions allows us to generate positives as well as hard-negatives with minimal additional overhead, e.g., we do not require an LLM for generating hard-negatives [9], nor do we need to generate synthetic images [41]. In Figure˜1, we show that CLIC consistently improves both on SugarCrepe++ and on image-to-text and text-to-image retrieval benchmarks compared to the pre-trained CLIP-ViT-B/32 model [45], whereas existing methods do not achieve consistent gains. To summarize, our contributions are as follows.
@@ 2 Methods and benchmarks for improving compositionality
Recent studies on VLMs have revealed significant limitations in their compositional abilities [56, 18, 52, 36, 42] via specially designed benchmarks. Throughout these works, several types of image-text models have been shown to lack compositional reasoning, ranging from those using just contrastive loss (CLIP [45], LaCLIP [12], SigLiP [57]) to the ones using different losses or architectures like CoCa [55], FLAVA [48], CLIPA [32], BLIP [16]. In this section, we introduce our notation for standard contrastive learning and provide details on methods to improve compositionality and the associated benchmarks.
@@ 2.2 Compositionality enhancing methods
Dense and Aligned Captions (DAC) [9] uses the same method to create hard negatives, but first generates higher quality captions by passing the image through BLIP-2 [25]. In DAC-LLM, the enhanced captions are used as input to GPT-Neo-2.7B [14] for detailed and dense captions, and in DAC-SAM, segments of the original image, created using SAM [21], are captioned via BLIP-2. The model is then fine-tuned using a combination of (Eq.˜1), (Eq.˜3) and MIL-loss adapted from [38], which uses multiple negatives for each image.
TripletCLIP [41] uses rewritten captions from LaCLIP [12] and utilizes an LLM to generate hard-negatives by changing concepts such as objects, attributes, and relationships that change the meaning of the original caption while slightly distorting it. Afterward, a synthetic image is generated for the hard-negative captions using a text-to-image diffusion model. Finally, training is done by using a sum of image-text and text-image based from Eq.˜2.
@@ 2.3 Compositionality benchmarks
Recently, SugarCrepe++ [11] extended this idea by introducing , an additional positive caption that is semantically similar to , but lexically different from both and . We note that SugarCrepe++ uses the additional positive caption , for the same images, original positive captions, and plausible negative captions from SugarCrepe. To answer correctly, the model needs to assign a higher score to and compared to the score assigned to . This makes SugarCrepe++ harder and more meaningful than SugarCrepe, as the model can no longer rely solely on lexical differences but has to instead understand the semantic differences between the captions. Therefore, we predominantly focus on SugarCrepe++ in this work.
@@ 3 Compositionally-aware learning in CLIP
Large-scale web-crawled datasets are known to be noisy, with many captions not reflecting their corresponding images [27, 3, 59]. This can be detrimental to compositional learning, which relies on minor modifications that alter semantic meaning. To improve the compositionality, we opt for captions that better align with the images and describe the images in detail. For this, we use approximately sample subsets taken from different data sources, like, RedCaps and CC12M from PixelProse [50], and our CogVLM [53] recaptioned Laion. Details about the datasets can be found in Appendix˜B. We train several versions of CLIC using each dataset to demonstrate the generalizability of our approach across data sources. Additionally, since some baselines are trained using MS-COCO, for a fair comparison, we train another version on this dataset, the details of which are in Section˜4.
@@ 3.1 Generation of positives and hard-negatives: concatenation of images and captions
To overcome the “bag of words" limitation, we aim for the model to distinguish between captions that are lexically similar but semantically different. In order to achieve this, we construct hard negatives that retain the same words in the caption but no longer correctly describe the image. For efficiency and diversity, we create these examples by concatenating pairs of images and swapping words across their two captions, ensuring that the resulting caption does not represent the concatenated images. To help the model learn that different lexical or syntactic sentences can represent the same image, we create multiple positives. For this, we use dense captions that provide varied descriptions of each image. The concatenation enables the number of combinations to grow quadratically with the dataset size, allowing CLIC to generalize better than other methods, see discussion in Section˜B.3. We detail our data preparation strategy based on concatenating random image-caption pairs in the following section. The process is illustrated in Figure 2 and described in Algorithm 1 in Appendix C. See Section˜B.4 for a discussion on a non-random approach.
@@
: We then sample (without repetition) two additional sentences from each caption and concatenate them in a random order , and for random indices . This step reinforces the model’s ability to distinguish incorrect captions from correct ones, even when the incorrect caption presents a general description of the image while the correct caption describes only a specific part. This is different from previous works, such as DAC [9], where multiple positives exist, but each caption faces its own hard-negative. An ablation on the number of positives can be found in Table˜21.
: Next, we leverage the spaCy package [17] to generate hard-negatives. We decompose the first two sentences into individual words and identify their respective linguistic categories using spaCy. A table detailing these categories, along with examples, is provided in Table 7 in Appendix B. To construct hard-negatives, we randomly select a linguistic category that is common to both sentences and choose two words from that category – one from each image. If no common category exists, we randomly select a word from each image and swap them. Certain categories, such as punctuation, are excluded from this process (see Appendix B.2 for details). To prevent over-fitting to specific benchmarks, we do not impose constraints on the selection of particular categories or word positions. As a result, the generated sentence no longer accurately represents the concatenated image unless, by chance, the swapped words share the same meaning.
@@ 3.2 Training of CLIC
Hard-negative loss: We compute a hard-negative loss between each of the positive captions and the negative caption of the same image, similar to Eq.˜3,
@@
where , , and are hyperparameters controlling the contribution of each loss term. To prevent the model from deviating too much from the pretrained CLIP when training only on concatenated images, we employ the standard clip training every second iteration, using a single image, the first sentence, and the standard clip loss as in Eq.˜1.
@@ 3.3 Differences between CLIC and other methods
Other approaches for generating hard text negatives differ from ours in computational cost and methodology. In DAC [9] and SVLC [10], hard-negatives tailored explicitly for specific compositionality tasks (e.g., replacing specific categories like colors, actions, etc.) are either found through look-up tables or created with computationally expensive LLM-based generation. Similarly, TripletCLIP uses an LLM to generate hard-negatives; for details, see Section˜2.2. Closest to our work is NegCLIP, which also uses word swapping. However, NegCLIP focuses on specific word classes such as adjectives, adverbs, verbs, and nouns as described in Appendix˜D. We allow a swap of any two random words as long as they are from the same class, e.g., prepositions, numbers, and pronouns, as seen in Section˜B.2. Our approach offers two main advantages: cheaper computational cost compared to LLM-based generation and a design that doesn’t target specific benchmarks, reducing the risk of over-fitting on existing benchmarks. Other methods also add hard-negative images, further increasing overall computational complexity. NegCLIP and CoN-CLIP search for hard images to add to the batch, and TripletCLIP generates hard images using a diffusion model. In addition, none of the methods use a combination of image pairs and corresponding text descriptions as in our method.
@@ 4.1 Experimental setup
Dataset. For CLIC, we fine-tune models either with CogVLM [53] recaptioned Laion images, or PixelProse [50] recaptioned images from RedCaps and CC12M (see Appendix˜B for details). We stress that this approach keeps our training data independent of MS-COCO, which is the basis for benchmarks like SugarCrepe, SugarCrepe++, and retrieval evaluations. By that, we ensure CLIC always functions as a zero-shot model. However, to fairly compare with the ViT-B/32 model of NegCLIP, which is trained on MS-COCO, we also fine-tune one model on the MS-COCO [29] dataset. In our experiments, we denote each model by the respective dataset it was trained on, i.e., CLIC-COCO, CLIC-LAION, CLIC-RedCaps, or CLIC-CC12M.
Models. To be comparable with previous compositionality enhancing methods, we base the majority of our experiments and the ablations on ViT-B/32 models pre-trained by OpenAI [45]. However, to highlight the adaptability and scalability of our approach, we also fine-tune other commonly used pre-trained architectures, such as ViT-B/16 and ViT-L/14 from [45]. Up to our knowledge, among methods improving compositionality, only CoN-CLIP [49] provides checkpoints for these architectures. Additionally, we fine-tune more recent ViT-L equivalent VLMs, including CLIPS and CLIPA to show that our method can enhance compositionality even in models that already possess compositional abilities and are pre-trained on varying datasets such as Re-DataComp (CLIPS), WIT (CLIP) and DataComp (CLIPA). More model specific details can be found in Section˜D.1.
Training details. In all of our experiments for CLIC, we keep the vision encoder frozen and fine-tune only the text encoder at an image resolution of . An ablation showing that this works better than fine-tuning the entire model can be found in Table˜20 in the appendix. Only for experiments associated with LLaVA [32], we fine-tune the complete model at a higher resolution of . We defer other training details like LR, loss parameters in Eq.˜7, etc, and a discussion on our NegCLIP baselines (denoted everywhere with a †), used whenever no checkpoints are available to Appendix˜D.
Evaluation. We evaluated the models on the compositionality benchmarks SugarCrepe++ [11], WinoGround [52], and SugarCrepe [18]. Additionally, we assess the effects of fine-tuning on downstream tasks such as classification on ImageNet [6] and on ZS-10, which computes the average score across ten standard classification datasets, detailed in Section˜D.2. In addition, we report image () and text retrieval () Recall@5 scores on MS-COCO (Val2017) [29] and Flickr30k [43]. We show Recall@1 numbers for retrieval and detailed SugarCrepe and SugarCrepe++ numbers for all the models in Appendix˜E.
@@ 4.2 Comparing to other methods
Given that most publicly available checkpoints of compositionality methods use ViT-B/32 based CLIP, we first focus on this architecture. From Table˜1, we observe that most methods perform worse than the pre-trained CLIP on SugarCrepe++. Notably, only NegCLIP, CoN-CLIP, and the CLIC models improve or are on par with the pre-trained baseline. Among these, CLIC-COCO is better than NegCLIP and CoN-CLIP (all three were trained on MS-COCO), while CLIC-RedCaps has lower TOT with a higher ITT (10.2% gain on pre-trained model across replace and swap). We also evaluate all models on the WinoGround benchmark, where, apart from CLIC, no other method improves on the pre-trained model, highlighting the effectiveness of CLIC. Moreover, CLIC improves in compositionality regardless of the fine-tuning dataset, showing that CLIC works for different sources of captions and images used for fine-tuning. An interesting finding in Table˜1 is that many models perform well on ITT in SugarCrepe but struggle on SugarCrepe++ ITT, an extended version including lexically different but semantically similar positives. This stark difference for methods like DAC (achieves 89.4% on SugarCrepe replace ITT but only 53.7% on SugarCrepe++) suggests these models detect lexical changes rather than truly learning compositional relations. This likely stems from their over reliance on hard-negatives tuned for previous benchmarks (see Section˜D.4).
Downstream tasks. We evaluate models on standard retrieval benchmarks (MS-COCO, Flickr30k) and classification tasks (ImageNet, ZS-10) in Table˜2. While TripletCLIP performs well on SugarCrepe++ and WinoGround, it exhibits significant degradation on ImageNet and ZS-10. In contrast, our CLIC models show minimal performance degradation on classification tasks, with CLIC-CC12M achieving the highest text retrieval performance among models not trained on MS-COCO while also improving image retrieval over the pre-trained baseline.
@@ 4.3 Ablation Study for CLIC
| Downstream Evaluations | Sugarcrepe++ | Sugarcrepe | ||||||||||
| IMNET | ZS-10 | Text | Image | ITT | TOT | ITT | TOT | ITT | ITT | ITT | ||
| Single image Baselines (Non-Concat) | ||||||||||||
| Ablating CLIC (Concat) | ||||||||||||
| +CLIP iterate (CLIC) | 61.7 | 61.0 | 75.9 | 60.0 | 75.6 | 60.1 | 61.1 | 27.9 | 84.5 | 84.0 | 73.7 | |
In Table˜4 we ablate the components of CLIC in steps (C1-C5) and present the analogous non-concatenated, single image version of CLIC (B1-B4). The CLIC baseline (C1) is fine-tuning on the concatenated images with caption and in (C2) we add our single hard negative. These two steps are sufficient to improve on SugarCrepe but not on SugarCrepe++. This suggests that the hard-negative distractor helps the model to detect lexical changes, but not to distinguish between semantically similar sentences and ones with different meaning. Adding multiple positives (C3) and the uni-modal loss (C4) improves significantly on SugarCrepe++ while performance on SugarCrepe slightly reduces. Thus, adding diverse positives and enforcing invariance of semantically identical captions lets the model identify semantically similar captions while being sensitive to lexically similar but semantically dissimilar captions. In the last step (C5) we add the iterates with single images for standard contrastive loss, which further improves compositionality and downstream task performance by eliminating the potential concatenated image biases. The non-concatenated, single-image variant of CLIC shows similar but smaller improvements in compositionality while being significantly worse in retrieval. Here, the NegCLIP† variant (similar to B2, but employing (2) in place of (3)) was used to compare CLIC to NegCLIP on a dataset other than MS-COCO, in order to avoid potential test data leakage from MS-COCO during training, as discussed in Section˜2.3.